Find the general/particular solution of the following Differential Equations

(Exact D.E)


4.) (dx/y) - x (dy/y^2)=0


1
Expert's answer
2021-09-07T18:29:04-0400

The equation that we have to solve is:


"\frac{1}{y}{dx}-\frac{x}{y^2}{dy}=0 \iff M{dx}+N{dy}=0"


Then we confirm this is an exact differential equation when we confirm that "\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}":

"\\M=\dfrac{1}{y};N=-\dfrac{x}{y^2};\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}=-\dfrac{1}{y^2}"


Then we proceed to solve for F(x,y):

"\\ F(x,y)=\int M dx=\dfrac{1}{y}\int {dx}=\dfrac{x}{y}+g(y)
\\ \frac{\partial F(x,y)}{\partial y}=-\dfrac{x}{y^2}+g\,'(y)=N
\\ N=-\dfrac{x}{y^2} \implies g\,'(y)=0 \iff g(y)=C=constant
\\ \implies \text{in conclusion } F(x,y)=\dfrac{x}{y}+C"